Circular No. 6735

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970828 S. C. Odewahn, S. G. Djorgovski, and S. R. Kulkarni, California In-

stitute of Technology; and D. A. Frail, National Radio Astronomy Observatory, report: “We examined CCD images of the field of GRB 970828 obtained at the Palomar 5-m Hale telescope by T. Herter, F. Fang, X. Xu, and O. Pevunova on Aug. 29 UT, and by C. Steidel, K. Adelberger, and M. Kellog on Aug. 30, 31, and Sept. 1. Our preliminary analyis shows, within the ASCA/RXTE/IPN error box, no objects brighter than magnitude R ∼ 24.5 whose brightness changed by > 0.2 mag/day in a statistically significant manner. This places a severe limit on the peak brightness of any possible optical transient associated with this burst, about a factor of 100 lower than the optical transient associated with GRB 970508.”

K. Z. Stanek, M. R. Garcia, and M. Krockenberger, Harvard-Smithson-ian Center for Astrophysics, communicate: “Analysis of images centered on the error box of the x-ray counterpart to GRB 970828 detected by RXTE (IAUC 6726, 6728) reveals no significantly-variable objects to a magnitude limit of R = 22. The variability for the 460 point sources in the field with 15.5 < R < 22 is rms < 0.3 mag. The field contains numerous well-resolved galaxies. The images were obtained with the 1.2-m telescope of the F. L. Whipple Observatory on Aug. 30.18, 30.19, Sept. 2.26, 2.27, 3.23, 3.24, 4.12, and 4.14 UT in the Cousins R filter (15-min exposures). The images cover an area 11′ × 11′ and are available as FITS files by anonymous ftp at cfa-

ftp.harvard.edu, in pub/kstanek/GRB970828, or through the WWW at
http://cfa-www.harvard.edu/~kstanek/. The faint (R = 24.5 ± 0.5)

optical object coincident with the variable radio source (IAUC 6730) is undetected on all of our individual and daily-summed images, which reach a detection threshold of R ∼ 22. Note that the position of this radio source is outside the 30′′ revised ASCA error radius (IAUC 6732).”

USNO 1425.09823278 Visual magnitude estimate (cf. IAUC 6731) by P. Schmeer, Bisch-

misheim, Germany: Sept. 5.021 UT, 13.5.

COMET C/1996 J1 (EVANS-DRINKWATER) Visual m_{1} estimates by R. J. Bouma, Groningen, The Netherlands

(0.25-m reflector): Aug. 5.98 UT, 13.7; 6.99, 13.6; Sept. 1.07, 13.7.

1997 September 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green